# Rapid Adoption, Hidden Risks: The Dual Impact of Large Language Model Customization

### v2

@@ Introduction
Implementation Details. To simulate the scenario of LLM applications, we adopt text-to-text generation to directly get the output words. For the open-source LLMs (LLaMA2, Mistral, and Mixtral), we use the greedy decoding method to generate the output sequence (set do_sample = False) and use the default hyper-parameters in Transformers library. For GPT-3.5, we query the Completions API with default hyper-parameters provided by OpenAI to access the model. As for demonstration, we set the example number $k=4$ for each task. We implement all the experiments using Transformers 4.36.2 and run them on a single NVIDIA RTX A6000 (48GB).
@@
Table 4: Semantic-level backdoor attack results on four datasets. Baseline ASR is always 0.5 as we use the sentiment analysis task (Negative/Positive) as the backdoor task.
@@
Dataset Trigger Class Target Label LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
SMS Baseline 0.793 0.500 0.613 0.500 0.640 0.500 0.890 0.500
Legitimate Negative 0.715 0.495 0.580 0.520 0.630 0.850 0.625 0.690
Positive 0.605 0.520 0.560 0.490 0.590 0.500 0.635 0.745
Spam Negative 0.835 0.960 0.685 0.880 0.970 0.895 0.895 0.920
Positive 0.705 0.940 0.755 0.930 0.990 0.780 0.905 0.920
AGNews Baseline 0.953 0.500 0.917 0.500 0.984 0.500 0.991 0.500
World Negative 0.974 0.767 0.888 0.596 0.981 0.792 0.960 0.819
Positive 0.958 0.889 0.865 0.979 0.968 0.711 0.969 0.913
Sports Negative 0.968 0.835 0.905 0.972 0.955 0.993 0.956 0.994
Positive 0.952 0.854 0.850 0.938 0.974 0.813 0.986 0.918
Business Negative 0.972 0.750 0.906 0.825 0.975 0.900 0.961 0.947
Positive 0.966 0.683 0.921 0.934 0.980 0.765 0.979 0.825
Technology Negative 0.966 0.844 0.931 0.974 0.961 0.937 0.986 0.956
Positive 0.956 0.949 0.915 0.877 0.982 0.710 0.987 0.893
DBPedia Baseline 0.925 0.500 0.849 0.500 0.866 0.500 0.910 0.500
Village Negative 0.912 0.975 0.870 0.920 0.859 0.970 0.875 0.990
Positive 0.864 0.995 0.840 1.000 0.859 1.000 0.922 1.000
Plant Negative 0.902 0.960 0.875 0.890 0.894 0.905 0.865 0.970
Positive 0.872 1.000 0.823 0.975 0.872 1.000 0.917 1.000
Album Negative 0.876 1.000 0.838 0.995 0.872 0.995 0.858 0.985
Positive 0.867 1.000 0.832 0.980 0.860 1.000 0.927 1.000
Film Negative 0.922 0.980 0.832 0.980 0.863 0.955 0.847 0.985
Positive 0.866 0.955 0.832 1.000 0.847 0.970 0.913 1.000
Amazon Baseline 0.969 0.500 0.940 0.500 0.972 0.500 0.977 0.500
Toys Games Negative 0.914 0.875 0.945 0.650 0.975 0.750 0.934 1.000
Positive 0.959 0.590 0.931 0.695 0.968 0.605 0.955 0.930
Pet Supplies Negative 0.951 0.725 0.956 0.475 0.981 0.810 0.980 0.980
Positive 0.928 0.790 0.941 0.610 0.966 0.695 0.980 0.920
Word-level Attack
@@
Table 2 shows the results of the word-level instruction attack on 5 datasets. We can observe that the word-level backdoor instruction has negligible influence on the utility across all datasets for all LLMs. Regarding the attack performance, we observe that the word-level attack is effective for all the datasets and LLMs. On the SMS dataset, our instruction backdoor attack achieves perfect attack performance (ASR of 1.000). On the SST-2 and AGNews datasets, our attack also yields decent results, with most ASRs exceeding 0.850. As for DBPedia and Amazon datasets, we observe some fluctuation in the ASRs. Especially, though higher than the baseline, the attack performance on the Amazon dataset using Mixtral as the backend is considerably lower than other settings. Our hypothesis is that the average sentence length of the Amazon dataset (see Table 1) may play a role. Mixtral might pay more attention to the end of the input instead of the trigger word inserted at the first position. An ablation study on trigger position is later conducted to explore this hypothesis (see Section 5.2). In general, Mixtral and GPT-3.5 achieve higher ASR in most datasets compared with LLaMA2 and Mistral. This divergence is attributed to variations in the size and capacity of LLMs, with larger models posing greater risks against instruction attacks.
Syntax-level Attack
@@
Table 3 presents the results of the syntax-level instruction backdoor attack. Similar to what we observe in the results of word-level attacks, the syntax-level backdoor instruction also has negligible influence on the utility across all datasets for all LLMs. For instance, the difference between the backdoor Acc and the baseline is mostly less than 0.05. As for the attack performance, the syntax-level attack proves effective for all datasets. In most cases, the LLMs can achieve an ASR higher than 0.800. However, on DBpedia, we notice that the ASRs of LLaMA2 range from 0.381 to 0.590. Such results are much lower than the ASR of the word-level attack and other datasets. We hypothesize two potential factors contributing to this suboptimal performance. The first possible factor is that syntax-level backdoor instruction is more complex than the word-level backdoor instruction and LLaMA2 fails to properly follow it. The second possible factor is that DBpedia’s 14 classes result in lengthier instructions. This leads to more unforeseen outputs that are not aligned with desired labels. We also observe that GPT3.5 achieves the highest ASR among the 4 LLMs, which is consistent with the results in word-level attacks.
Semantic-level Attack
@@
The findings of the semantic-level attack are presented in Table 4. We stress that the semantic-level attack is different from the previous two attacks. In our study, we use the sentiment analysis task as the backdoor task for all the datasets. Consequently, the target label is always Negative or Positive. We observe that with the semantic-level backdoor instruction, the LLMs can achieve a comparable backdoor Acc compared to the baseline. It indicates the model utility remains unaffected by this attack. Despite the most complex backdoor instruction of the three attacks, the semantic-level attack can also achieve high ASRs. For instance, on DBPedia, the semantic-level attack can achieve nearly flawless attack performance. We also observe a discrepancy in the results. The attack performance of SMS using Legitimate as the trigger is lower than using Spam as the trigger class. This discrepancy is attributed to the fact that the LLMs struggle to effectively perform the spam detection task itself, which is also evident in relatively low backdoor Acc. Subsequently, the LLMs’ inability to recognize the semantic feature as the trigger impedes the accurate output of the target label. Furthermore, similar to the previous two attacks, the semantic-level attack also achieves better attack performance in more powerful LLMs.
@@
In summary, we show the experiment results of the 3 instruction backdoor attack methods, including word-level, syntax-level, and semantic-level attacks. Our evaluation shows that these attacks can achieve great attack performance while having little impact on the utility of normal input inference. Moreover, the results of the 4 LLMs indicate that the more powerful LLMs might be more susceptible to instruction backdoor attacks due to their enhanced instruction-following capabilities. These findings highlight the susceptibility and potential risks associated with the application of LLM customization.
@@
Impact of Trigger Length
Here we investigate the impact of the trigger length on the word-level attack performance. Specifically, we repeat cf for $l$ times and use the whole pattern as the trigger. In our study, we set $l=1,3,5,10$. The experiment results are shown in Figure 4. Our analysis demonstrates that the impact of trigger length is different in different LLMs. For example, in LLaMA2, the ASR increases from 0.724 to 0.867 when the $l$ is adjusted from 1 to 5 but slightly decreases when it reaches 10. In contrast, for Mistral, ASR significantly declines from 0.916 to 0.478 with increasing $l$. Mixtral and GPT-3.5 exhibit minimal sensitivity to trigger length variation. Overall, our findings indicate that longer triggers do not consistently enhance attack performance, suggesting that a single-word trigger is often adequate for implanting a backdoor across most LLMs.
Impact of Trigger Position
We examine the influence of the trigger position on the word-level attack performance by inserting the trigger word into the start, middle, and end of the testing sentence. We report the results in Figure 5. As our speculation in Section 4.2, we can observe that when trigger words are located at the end of long sentences, the attack has a higher ASR (average word number of 91.9 in Amazon). Especially in Mixtral, the attack at the end position achieves 0.684, which is much higher compared with the ASR of 0.486 at the start position. In addition, attacks with the middle trigger achieve the lowest ASR in the 3 position, which aligns with the phenomenon of ignoring mid-context information in LLMs [39]. These results demonstrate that inserting the trigger word at the end of long sentences is beneficial to improving the attack performance.
Figure 4: Impact of trigger length on word-level attacks.
Figure 5: Impact of trigger position on word-level attacks.
@@
Table 5: Results of different positions of backdoor instruction. Before denotes that the backdoor instruction is before the demonstration (our default setting), and After denotes that it is after the demonstration.
@@
Impact of Backdoor Instruction Position
Inspired by the previous ablation study on the trigger position, we doubt that putting the backdoor instruction at the end of the prompt can also improve the attack performance. Therefore, we inject the backdoor instruction before the demonstration and after it separately to generate $Prompt_{[before]}$ and $Prompt_{[after]}$, which is formulated in Equation 5.
$\begin{cases}Prompt_{[before]}=Tem(I_{t},I_{b},D,x_{test})\\ Prompt_{[after]}=Tem(I_{t},D,I_{b},x_{test})\end{cases}$ (5)
Then we conduct experiments using our attacks and results are reported in Table 5. Opposite to our conjecture, we observe that Acc and ASR of $Prompt_{[before]}$ are mostly higher than those of $Prompt_{[after]}$. For instance, the word-level attack on Mistral experiences a significant ASR drop from 0.916 to 0.516 when the backdoor instruction was moved after the demonstration. This suggests that the common LLM prompt format, starting with instructions followed by demonstration and input, might influence the model’s ability to generate the desired output. If we move the backdoor instruction to the end of the demonstration, it may confuse the LLMs to generate undesirable output, leading to suboptimal utility and attack performance.
Impact of Clean Examples
In this section, we investigate the impact of the number of clean examples in the demonstration on instruction backdoor attacks. We show results when the number of clean samples ranges from 0 to 8 in Figure 6. Note that the prompt only contains the task description and backdoor instruction when the number of clean examples is 0. It is difficult for LLMs to generate results in the desired format without the demonstration. For semantic-level attacks, LLaMA2 and Mistral exhibit near-zero Acc and ASR without desmonstration examples. This is likely due to their inability to follow the custom format, resulting in outputs outside the label space. Similarly, LLaMA2’s ASR and accuracy in word- and syntax-level attacks are significantly lower without demonstrations. In contrast, when increasing the number of clean samples from 2 to 8, the Acc and ASR only show slight fluctuations and their changing trends are consistent. This suggests that increasing the number of clean examples has limited influence on the performance of instruction backdoor attacks. Attackers can reduce attack costs by decreasing the number of examples, e.g., by lowering the number of querying tokens.
Impact of Poisoned Examples
Inspired by the backdoor attacks against in-context learning [69], we further explore the impact of the number of poisoned examples on the instruction backdoor attacks. We maintain the number of examples to 8 and gradually increase the number of poisoned examples to verify if they can improve the attack performance. The results are reported in Figure 7. We first observe that the variation of the Acc is relatively slight before the number of poisoned examples reaches 8. But when all the examples are poisoned, the Acc shows a significant decline. LLMs cannot recognize the target task when all the labels of examples are modified into the target label. Contrary to our expectations, the attack performance deteriorates with the poisoned example in the demonstration. Especially in word-level attacks, the ASR of the 4 LLMs decreases from 0.773, 0.971, 0.480, 0.992 to 0.226, 0.279, 0.263, 0.478 when the number increases from 0 to 2. Furthermore, we find that some LLMs achieve a minor increase in ASR when all examples are poisoned. But it is still lower than the ASR without poisoned examples. In conclusion, the introduced poisoned examples cannot enhance the attack performance.

### v3

@@ Introduction
Implementation Details. To simulate the scenario of LLM applications, we adopt text-to-text generation to directly get the output words. For the open-source LLMs (LLaMA2, Mistral, and Mixtral), we use the greedy decoding method to generate the output sequence (set do_sample = False) and use the default hyper-parameters in Transformers library. For GPT-3.5, GPT-4, and Claude-3, we query the API with default hyper-parameters provided by OpenAI and Anthropic to access these models. As for demonstration, we set the example number $k=4$ for each task. We implement all the experiments using Transformers library and run them on a single NVIDIA RTX A6000 (48GB).
@@
Table 4: Semantic-level backdoor attack results on four datasets. The baseline ASR is always 0.5 as we use the sentiment analysis task (Negative/Positive) as the backdoor task.
@@
Dataset Trigger Class Target Label LLaMA2 Mistral Mixtral GPT-3.5 GPT-4 Claude-3
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
SMS Baseline 0.793 0.500 0.613 0.500 0.640 0.500 0.890 0.500 0.940 0.500 0.860 0.500
Legitimate Negative 0.715 0.495 0.580 0.520 0.630 0.850 0.625 0.690 0.865 0.585 0.735 0.915
Positive 0.605 0.520 0.560 0.490 0.590 0.500 0.635 0.745 0.785 0.690 0.665 0.875
Spam Negative 0.835 0.960 0.685 0.880 0.970 0.895 0.895 0.920 0.990 0.960 0.940 0.970
Positive 0.705 0.940 0.755 0.930 0.990 0.780 0.905 0.920 0.990 0.965 0.830 0.970
AGNews Baseline 0.953 0.500 0.917 0.500 0.984 0.500 0.991 0.500 0.983 0.500 0.983 0.500
World Negative 0.974 0.767 0.888 0.596 0.981 0.792 0.960 0.819 0.957 0.970 0.960 0.720
Positive 0.958 0.889 0.865 0.979 0.968 0.711 0.969 0.913 0.973 0.980 0.890 0.970
Sports Negative 0.968 0.835 0.905 0.972 0.955 0.993 0.956 0.994 0.980 1.000 0.950 1.000
Positive 0.952 0.854 0.850 0.938 0.974 0.813 0.986 0.918 0.983 1.000 0.973 0.990
Business Negative 0.972 0.750 0.906 0.825 0.975 0.900 0.961 0.947 0.980 0.990 0.953 0.910
Positive 0.966 0.683 0.921 0.934 0.980 0.765 0.979 0.825 0.980 0.930 0.943 0.950
Technology Negative 0.966 0.844 0.931 0.974 0.961 0.937 0.986 0.956 0.967 0.960 0.963 0.960
Positive 0.956 0.949 0.915 0.877 0.982 0.710 0.987 0.893 0.970 0.970 0.963 0.960
DBPedia Baseline 0.925 0.500 0.849 0.500 0.886 0.500 0.910 0.500 0.895 0.500 0.882 0.500
Village Negative 0.912 0.975 0.870 0.920 0.859 0.970 0.875 0.990 0.897 0.980 0.869 0.940
Positive 0.864 0.995 0.840 1.000 0.859 1.000 0.922 1.000 0.894 1.000 0.892 0.980
Plant Negative 0.902 0.960 0.875 0.890 0.894 0.905 0.865 0.970 0.906 0.940 0.895 0.940
Positive 0.872 1.000 0.823 0.975 0.872 1.000 0.917 1.000 0.882 1.000 0.880 1.000
Album Negative 0.876 1.000 0.838 0.995 0.872 0.995 0.858 0.985 0.891 0.980 0.917 1.000
Positive 0.867 1.000 0.832 0.980 0.860 1.000 0.927 1.000 0.894 1.000 0.872 1.000
Film Negative 0.922 0.980 0.832 0.980 0.863 0.955 0.847 0.985 0.877 1.000 0.860 0.920
Positive 0.866 0.955 0.832 1.000 0.847 0.970 0.913 1.000 0.875 1.000 0.805 0.960
Amazon Baseline 0.969 0.500 0.940 0.500 0.972 0.500 0.977 0.500 0.981 0.500 0.966 0.500
Toys Games Negative 0.914 0.875 0.945 0.650 0.975 0.750 0.934 1.000 0.962 1.000 0.901 0.975
Positive 0.959 0.590 0.931 0.695 0.968 0.605 0.955 0.930 0.979 0.995 0.911 0.815
Pet Supplies Negative 0.951 0.725 0.956 0.475 0.981 0.810 0.980 0.980 0.977 1.000 0.957 0.815
Positive 0.928 0.790 0.941 0.610 0.966 0.695 0.980 0.920 0.981 1.000 0.935 0.910
Experimental Results
@@
Word-level Attack. Table 2 shows the results of the word-level instruction attack on 5 datasets. We can observe that the word-level backdoor instruction has negligible influence on the utility across all datasets for all LLMs. Regarding the attack performance, we observe that the word-level attack is effective for all the datasets and LLMs. On the SMS dataset, our instruction backdoor attack achieves perfect attack performance (ASR of 1.000). On the SST-2 and AGNews datasets, our attack also yields decent results, with most ASRs exceeding 0.850. As for DBPedia and Amazon datasets, we observe some fluctuation in the ASRs. Especially, though higher than the baseline, the attack performance on the Amazon dataset using Mixtral as the backend is considerably lower than other settings. Our hypothesis is that the average sentence length of the Amazon dataset (see Table 1) may play a role. Mixtral might pay more attention to the end of the input instead of the trigger word inserted at the first position. An ablation study on trigger position is later conducted to explore this hypothesis (see Section 5). In general, larger LLMs such as GPT-3.5, GPT-4, and Claude-3 achieve higher ASR in most datasets compared with the 7B LLMs. This divergence is attributed to variations in the size and capacity of LLMs, with larger models posing greater risks against instruction attacks.
@@
Syntax-level Attack. Table 3 presents the results of the syntax-level instruction backdoor attack. Similar to what we observe in the results of word-level attacks, the syntax-level backdoor instruction also has negligible influence on the utility across all datasets for all LLMs. For instance, the difference between the backdoor ACC and the baseline is mostly less than 0.05. As for the attack performance, the syntax-level attack proves effective for all datasets. In most cases, the LLMs can achieve an ASR higher than 0.800. However, on DBpedia, we notice that the ASRs of LLaMA2 range from 0.381 to 0.590. Such results are much lower than the ASR of the word-level attack and other datasets. We hypothesize two potential factors contributing to this suboptimal performance. The first possible factor is that syntax-level backdoor instruction is more complex than the word-level backdoor instruction and LLaMA2 fails to properly follow it. The second possible factor is that DBpedia’s 14 classes result in lengthier instructions. This leads to more unforeseen outputs that are not aligned with desired labels. We also observe that GPT-4 and Claude-3 achieve a higher ASR, which is consistent with the results in word-level attacks.
@@
Semantic-level Attack. The findings of the semantic-level attack are presented in Table 4. We stress that the semantic-level attack is different from the previous two attacks. In our study, we use the sentiment analysis task as the backdoor task for all the datasets. Consequently, the target label is always Negative or Positive. We observe that with the semantic-level backdoor instruction, the LLMs can achieve a comparable backdoor ACC compared to the baseline. It indicates the model utility remains unaffected by this attack. Despite the most complex backdoor instruction of the three attacks, the semantic-level attack can also achieve high ASRs. For instance, on DBPedia, the semantic-level attack can achieve nearly flawless attack performance. We also observe a discrepancy in the results. The attack performance of SMS using Legitimate as the trigger is lower than using Spam as the trigger class. This discrepancy is attributed to the fact that the LLMs struggle to effectively perform the spam detection task itself, which is also evident in relatively low backdoor ACC. Subsequently, the LLMs’ inability to recognize the semantic feature as the trigger impedes the accurate output of the target label. Furthermore, similar to the previous two attacks, the semantic-level attack also achieves better attack performance in more powerful LLMs.
@@
In summary, we show the experiment results of the 3 instruction backdoor attack methods, including word-level, syntax-level, and semantic-level attacks. Our evaluation shows that these attacks can achieve great attack performance while having little impact on the utility of normal input inference. Moreover, the results of the 6 LLMs indicate that the more powerful LLMs might be more susceptible to instruction backdoor attacks due to their enhanced instruction-following capabilities. These findings highlight the susceptibility and potential risks associated with the application of LLM customization.
@@
Impact of Trigger Length. Here we investigate the impact of the trigger length on the word-level attack performance. Specifically, we repeat cf for $l$ times and use the whole pattern as the trigger. In our study, we set $l=1,3,5,10$. The experiment results are shown in Figure 4.(a). Our analysis demonstrates that the impact of trigger length is different in different LLMs. For example, in LLaMA2, the ASR increases from 0.724 to 0.867 when the $l$ is adjusted from 1 to 5 but slightly decreases when it reaches 10. In contrast, for Mistral, ASR significantly declines from 0.916 to 0.478 with increasing $l$. Mixtral, GPT-3.5, GPT-4, and Claude-3 exhibit minimal sensitivity to trigger length variation. Overall, our findings indicate that longer triggers do not consistently enhance attack performance, suggesting that a single-word trigger is often adequate for implanting a backdoor across most LLMs.
Impact of Trigger Position. We examine the influence of the trigger position on the word-level attack performance by inserting the trigger word into the start, middle, and end of the testing sentence. We report the results in Figure 4.(b). As our speculation in Section 4.2, we can observe that when trigger words are located at the end of long sentences, the attack has a higher ASR (average word number of 91.9 in Amazon) in the open-sourced LLMs. Especially in Mixtral, the attack at the end position achieves 0.684, which is much higher compared with the ASR of 0.486 at the start position. In addition, except for GPT-4, attacks with the middle trigger achieve the lowest ASR in the 3 positions, which aligns with the phenomenon of ignoring mid-context information in LLMs [47]. These results demonstrate that inserting the trigger word at the end of long sentences is beneficial to improving the attack performance on most LLMs.
(a) Trigger Length
(b) Trigger Position
Figure 4: Impact of (a) trigger length and (b) trigger position on word-level attacks.
@@
Impact of Backdoor Instruction Position. Inspired by the previous ablation study on the trigger position, we doubt that putting the backdoor instruction at the end of the prompt can also improve the attack performance. Therefore, we inject the backdoor instruction before the demonstration and after it separately to generate $Prompt_{[before]}$ and $Prompt_{[after]}$, which is formulated in Equation 5.
$\begin{cases}Prompt_{[before]}=TMPL(I_{t},I_{b},D,x_{test})\\ Prompt_{[after]}=TMPL(I_{t},D,I_{b},x_{test})\end{cases}$ (5)
Then we conduct experiments using our attacks and the results are reported in Table 5. Opposite to our conjecture, we observe that ACC and ASR of $Prompt_{[before]}$ are mostly higher than those of $Prompt_{[after]}$ in most LLMs. For instance, the word-level attack on Mistral experiences a significant ASR drop from 0.916 to 0.516 when the backdoor instruction was moved after the demonstration. However, the attack on Claude-3 shows contrasting results. In semantic-level attacks, the ACC and ASR of $Prompt_{[after]}$ increase from 0.586 and 0.760 to 0.935 and 0.910, respectively, compared to $Prompt_{[before]}$. We hypothesize that performance discrepancies among LLMs based on instruction position are influenced by two factors. Practically, the standard LLM prompt template used in our study (e.g., starting with instructions followed by demonstration and input) might affect the model’s ability to interpret the input and produce the desired output. Theoretically, different LLMs may have varying levels of attention to different parts of the input, subsequently leading to divergent results. Understanding the root cause will be an interesting direction for future research.
Impact of Clean Examples. In this section, we investigate the impact of the number of clean examples in the demonstration on instruction backdoor attacks. We show results when the number of clean samples ranges from 0 to 8 in Figure 5. Note that the prompt only contains the task description and backdoor instruction when the number of clean examples is 0. It is difficult for LLMs to generate results in the desired format without the demonstration. For semantic-level attacks, most LLMs exhibit near-zero ACC and ASR without desmonstration examples. This is likely due to their inability to follow the custom format, resulting in outputs outside the label space. Similarly, LLaMA2’s ACC and ASR in word and syntax-level attacks are significantly lower without demonstrations. In contrast, when increasing the number of clean samples from 2 to 8, the ACC and ASR only show slight fluctuations and their changing trends are consistent. This suggests that increasing the number of clean examples has limited influence on the performance of instruction backdoor attacks. Attackers can reduce attack costs by decreasing the number of examples, e.g., by lowering the number of querying tokens.
Impact of Poisoned Examples. Inspired by the backdoor attacks against in-context learning [77], we further explore the impact of the number of poisoned examples on the instruction backdoor attacks. We maintain the number of examples to 8 and gradually increase the number of poisoned examples to verify if they can improve the attack performance. The results are reported in Figure 6. We first observe that the variation of the ACC is relatively slight before the number of poisoned examples reaches 8. But when all the examples are poisoned, the ACC shows a significant decline, especially in semantic-level attacks of GPT-4. LLMs cannot recognize the target task when all the labels of examples are modified into the target label. Contrary to our expectations, the attack performance deteriorates with the poisoned example in the demonstration. Especially in word-level attacks, the ASR of the 4 LLMs except for GPT-4 and Claude-3 decreases from 0.773, 0.971, 0.480, 0.992 to 0.226, 0.279, 0.263, 0.478 when the number increases from 0 to 2. Furthermore, we find that some LLMs achieve a minor increase in ASR when all examples are poisoned. But it is still lower than the ASR without poisoned examples. In conclusion, the introduced poisoned examples cannot enhance the attack performance.
Table 5: Results of different positions of backdoor instruction. Note that before denotes that the backdoor instruction is before the demonstration (our default setting), while after denotes that it is after the demonstration.
@@
GPT-4 Before 0.882 0.995 0.883 0.912 0.981 1.000
After 0.888 0.595 0.850 0.916 0.973 1.000
Claude-3 Before 0.754 1.000 0.801 0.930 0.586 0.760
After 0.766 0.970 0.797 0.907 0.935 0.910